OpenStudy (anonymous):

Consider a triangle with a=3, b=7 and C=37 degree angle. Use the Law of Cosines to find small c (side). Round to nearest tenth.

OpenStudy (anonymous):

do you know the law of cosines ?

OpenStudy (anonymous):

 c^2=a^2+b^2-2ab*cos(c) =9+49-42cos(37) c=4.9

OpenStudy (anonymous):

Do you mean the c^2=a^2 +b^2-2abcos(c)

OpenStudy (anonymous):

oh wow, thank you. my computer seems to be getting these responses slow.

OpenStudy (anonymous):

$$3^2+7^2-2\times 21 \cos(37^\circ) = 24.4573 $$

OpenStudy (anonymous):

now take the square root to get 4.94543